### v1

@@ 3 Convergence of Model-based Policy Gradient
  • Gradient Descent: $K_{i+1}=K_{i}-\eta\nabla C(K_{i})$;
@@
where $i$ is the iteration index and $i\in\mathbb{Z}_{+}$. The Gauss-Newton method, when implemented with a step size $\eta=\frac{1}{2}$, becomes equivalent to the policy iteration algorithm, which is well-known for its quadratic convergence rate [13]. To implement the Gauss-Newton method, three terms, $(R+B^{\top}P_{K_{i}}B),\nabla C(K_{i}),\Sigma_{K_{i}}$ are necessary. In contrast, gradient descent only requires the gradient $\nabla C(K_{i})$, but exhibits only a linear convergence rate. The natural policy gradient method strikes a balance between the two, requiring less information than the Gauss-Newton method while providing faster convergence properties than the gradient descent method.
3.1 Gradient Descent
In this subsection, we present the theorem guaranteeing the convergence of the model-based gradient descent method, based on the properties introduced in Sections 2.1 and 2.2. {Theorem}[Gradient Descent with Adaptive Step Size] Suppose the initial $K_{0}\in\mathcal{S}$, and consider the gradient descent iteration
(14) $K_{i+1}=K_{i}-\eta_{i}\nabla C(K_{i}),\quad\forall i\in\mathbb{Z}_{+},$
where the step size satisfies $\eta_{i}\leq h_{\mathrm{GD}}(C(K_{i}))$, and $h_{\mathrm{GD}}(C(K_{i}))$ is defined in (62) in Appendix B.1. Then, the following relationship holds:
(15) $C(K_{i+1})-C(K^{*})\leq\left(1-\frac{2\eta_{i}\lambda_{1}{(R)}\lambda_{1}^{2}(\Sigma_{w})}{\lVert\Sigma_{K^{*}}\rVert}\right)(C(K_{i})-C(K^{*})).$
For $\eta_{i}=h_{\mathrm{GD}}(C(K_{i})),i\in\mathbb{Z}_{+}$ and given any accuracy gap $\epsilon>0$, if the number of iterations $N$ satisfies:
@@ 3.1 Gradient Descent
The proof of the Theorem 3.1 can be found in Appendix B.1. From Theorem 3.1, we see that the gradient descent method can achieve any desired accuracy $\epsilon$ within a finite number of iterations $N$ and converges linearly. According to Theorem 3.1, the step size is adaptive according to the cost $C(K_{i})$ at the current iterate $K_{i}$.
{Remark}
[Advantage of Adaptive Step Size over Fixed Step Size] In [10], a fixed step size is introduced. From Theorem 3.1, a fixed step size version can also be derived, where the fixed step size is given by $\eta=h_{\mathrm{GD}}(C(K_{0}))$. For both fixed and adaptive step sizes, we can prove that $C(K_{i+1})\leq C(K_{i})$ from Theorem 3.1. Given the same initial $K_{0}$ and together with the expression of $h_{\mathrm{GD}}(C(K))$, it follows that $\eta_{0}\leq\eta_{i}$ for all $i$. This leads to the following inequality:
@@
Thus, it can be concluded that using an adaptive step size improves the convergence rate compared to a fixed step size.
@@ 3.2 Natural Policy Gradient
(16) $\theta\leftarrow\theta-\eta G_{\theta}^{-1}\nabla C(\theta),$
@@
(18) $\hat{G}_{\bar{K}}:=\lim_{T\rightarrow+\infty}\frac{1}{T}\sum_{t=0}^{T}~{}\mathbb{E}\left[\nabla\mathop{log}\pi_{\bar{K}}(u_{t}|x_{t})\nabla\mathop{log}\pi_{\bar{K}}(u_{t}|x_{t})^{\top}\right].$
@@
(19) $K_{i+1}=K_{i}-\eta\nabla C(K_{i})\Sigma_{K_{i}}^{-1}$
The explicit derivation of (19) starting from (16) and using the particular policy parametrization (17) is given in Appendix B.2. We now present the theorem that establishes the convergence guarantee of the natural policy gradient method {Theorem}[Natural Policy Gradient with Adaptive Step Size] Suppose the initial $K_{0}\in\mathcal{S}$, and consider the natural policy gradient iteration
(20) $K_{i+1}=K_{i}-\eta_{i}\nabla C(K_{i})\Sigma_{K_{i}}^{-1},\quad\forall i\in\mathbb{Z}_{+},$
@@
(21) $\eta_{i}\leq\frac{1}{2\lVert R\rVert+\frac{2\lVert B\rVert^{2}C(K_{i})}{\lambda_{1}(\Sigma_{w})}}.$
@@
(22) $C(K_{i+1})-C(K^{*})\leq\left(1-\frac{2\eta_{i}\lambda_{1}{(R)}\lambda_{1}(\Sigma_{w})}{\lVert\Sigma_{K^{*}}\rVert}\right)(C(K_{i})-C(K^{*})).$
@@
(23) $\eta_{i}=\frac{1}{2\lVert R\rVert+\frac{2\lVert B\rVert^{2}C(K_{i})}{\lambda_{1}(\Sigma_{w})}},\quad\forall i\in\mathbb{Z}_{+},$
@@
The proof of Theorem 19 relies on Lemma 12 and follows the procedure outlined in [10, Lemma 15]. Theorem 19 establishes the global convergence properties of the natural policy gradient method, when an appropriate step size $\eta_{i}$ is selected. Similar to the observation in Remark 3.1, the introduction of adaptive step sizes improves the convergence rate, as the step sizes increase adaptively with the decreasing of the cost.
@@ 3.3 Gauss-Newton Method
We now present the theorem that provides the convergence guarantee for the Gauss-Newton method. {Theorem}[Gauss-Newton Method] Suppose the initial $K_{0}\in\mathcal{S}$, and consider the Gauss-Newton iteration
(24) $K_{i+1}=K_{i}-\eta(R+B^{\top}P_{K_{i}}B)^{-1}\nabla C(K_{i})\Sigma_{K_{i}}^{-1},\quad\forall i\in\mathbb{Z}_{+},$
@@
(25) $C(K_{i+1})-C(K^{*})\leq\left(1-\frac{2\eta\lambda_{1}(\Sigma_{w})}{\lVert\Sigma_{K}^{*}\rVert}\right)(C(K_{i})-C(K^{*})).$
@@
The proof of the Theorem 3.3 builds on the property shown in Lemma 12 and follows similar steps to the steps in [10, Lemma 14]. It is important to note that the choice of step size here is independent of the specific system parameters. To achieve the fastest convergence rate, one can always select the step size $\eta=\frac{1}{2}$, which is equivalent to the policy iteration algorithm.
@@ 3.4 Lack of robustness of model-based PG methods
(26) $x_{t+1}=\underbrace{\left[\begin{array}[]{ccc}1.01&0.01&0\\ 0.01&1.01&0.01\\ 0&0.01&1.01\end{array}\right]}_{A}x_{t}+\underbrace{\left[\begin{array}[]{ccc}1&0&0\\ 0&1&0\\ 0&0&1\end{array}\right]}_{B}u_{t}+w_{t}.$
@@
(27) $Q=0.001I_{3},\quad R=I_{3}$
The initial condition $\hat{K}_{0}$ is chosen as the optimal gain for the LQR problem with $(A,B,50Q,R)$. To model gradient estimation errors due to noisy data, we define the gradient estimate as follows:
(28) ^$\hat{\nabla}C(K):=\nabla C(K)+\Delta,$
where $\Delta\in\mathbb{R}^{3\times 3}$ is a matrix with entries $\Delta_{ij}$ sampled according to the distribution $\mathcal{N}(0,\sigma^{2})$, where the values of $\sigma\geq 0$ will be discussed later. The gradient descent algorithm updates the gain $\hat{K}_{i}$ as follows:
(29) $\hat{K}_{i+1}=\hat{K}_{i}-\eta\hat{\nabla}C(\hat{K}_{i}),\quad i\in\mathbb{Z}_{+}.$
@@
When the gradient is exactly known (i.e. $\sigma=0$, black dotted line), we can use theoretical bounds on the step size to guarantee convergence of the gradient descent algorithm. In this case, the converging behavior of $\eta=0.12$ to the optimal policy is shown as an example. When the gradient estimation is affected by a small amount of noise ($\sigma=0.03$, blue dashed line), the same step size yields convergence but to a suboptimal value of gain $K$. As the noise level increases ($\sigma=0.6$, green dash-dotted line), the algorithm starts showing non-converging behavior. In this case, reducing the step size addresses the issue, as shown by the red solid line where the step size is empirically decreased to $0.03$. It is worth observing that this reduction in step size results in a slower convergence rate.
These results exemplify that when PG algorithms are run using finite noisy data trajectories to compute gradients (or equivalently covariances), the convergence guarantees obtained in the model-based scenario no longer hold. Specifically, the algorithm might converge to suboptimal solutions or even diverge. It is also shown that, in the latter case, by properly modifying the algorithms’ parameters (such as the step size), converging behaviors can be achieved.
To characterize the fundamental limitations of such algorithms and improve their performance it is then paramount to: quantify the uncertainty on gradients and covariances due to the use of finite noisy data; and analyze how this uncertainty propagates in the policy optimization algorithm. To the best of the authors’ knowledge, there is no holistic analysis that captures all these aspects and provides guarantees for model-free PG algorithms in the presence of noisy trajectory data. The main technical contribution of this work is to bridge this gap by providing: probabilistic guarantees of convergence and suboptimality bounds; closed-form expressions for the algorithm’s parameters (e.g. step size, variance reduction) that guarantee converging behavior
@@ 4 Convergence of Model-free Policy Gradient
In this section, we analyze two variants of model-free policy gradient methods: gradient descent and natural policy gradient. Unlike model-based approaches, the system dynamics $(A,B)$ are unknown in the model-free setting. Instead of relying on system identification, we use zeroth-order optimization, where gradients and covariance estimates are estimated directly from trajectory data. In Section 4.1, we begin by analyzing estimation algorithms for two fundamental components utilized in the proposed methods: the cost function gradient and the covariance. We derive novel finite-sample error bounds that explicitly account for the effects of noisy data. Building on these quantitative bounds, Section 4.2 introduces a variance reduction technique to provably reduce the required number of rollouts. Leveraging these results, we establish probabilistic convergence guarantees for GD (Section 4.3) and NPG (Section 4.4), emphasizing the more general scenario with time-adaptive step sizes.
@@ 4.1 Model-free Gradient and Covariance Estimation
  for $k=1,...,n$ do
      1. Generate a sample gain matrix $\hat{K}_{k}=K+U_{k}$, where $U_{k}$ is drawn uniformly at random over matrices of compatible dimensions with Frobenius norm $r$;
@@ 2. Generate a sample initial state $x_{0}^{(k)}$ with $\lVert x_{0}^{(k)}\rVert\leq L_{0}$;
      3. Excite the closed-loop system with $u_{t}^{(k)}=\hat{K}_{k}x_{t}^{(k)}$ for $l$-steps starting from $x_{0}^{(k)}$, yielding the state sequence $\left\{x_{t}^{(k)}\right\}_{t=0}^{l-1}$ originating from (1b);
      4. Collect the empirical cost estimate $\hat{C}_{k}:=\frac{1}{l}\sum_{t=0}^{l-1}x_{t}^{(k)~{}\top}(Q+\hat{K}^{\top}_{k}R\hat{K}_{k})x_{t}^{(k)}$ and the empirical covariance matrix $\hat{\Sigma}_{i}=\frac{1}{l}\sum_{t=0}^{l-1}~{}x_{t}^{(k)}x_{t}^{(k)~{}\top}$;
@@
  return  Gradient estimate ^$\hat{\nabla}C(K):=\frac{1}{n}\sum^{n}_{k=1}\frac{n_{x}n_{u}}{r^{2}}\hat{C}_{k}U_{k}$ and covariance estimate $\hat{\Sigma}_{K}:=\frac{1}{n}\sum^{n}_{k=1}\hat{\Sigma}_{k}$.
We present two theorems to analyze the estimation errors of both ^$\hat{\nabla}C(K)$ and $\hat{\Sigma}_{K}$, obtained from Algorithm 1. {Theorem}[Error Bound of ^$\hat{\nabla}C(K)$] Given an arbitrary tolerance $\epsilon>0$, which can be decomposed as $\epsilon=\epsilon_{d}+\epsilon_{l}+\epsilon_{n}+\epsilon_{r}$, and an arbitrary probability $\delta\in(0,1)$, which can be decomposed as $\delta=(1-\delta_{d})(1-\delta_{n})\delta_{x}$, for a given $K\in\mathcal{S}$, the estimated gradient ^$\hat{\nabla}C(K)$ from Algorithm 1 enjoys the following bound:
(30) {^$\mathbb{P}\left\{\lVert\hat{\nabla}C(K)-{\nabla}C(K)\rVert\leq\epsilon\right\}\geq\delta,$
@@
(31a) $r\leq r_{\max}(C(K),\epsilon_{r}),$ $\quad l\geq l_{\min}(C(K),\epsilon_{l}),$
(31b) $n\geq\max(N_{1}(C(K),\epsilon_{r},\epsilon_{l},\epsilon_{n},\delta_{n},\delta_{d}),$ $N_{2}(C(K),\epsilon_{r},\epsilon_{l},\epsilon_{n},\epsilon_{d},\delta_{x},\delta_{n},\delta_{d})),$
where the detailed expressions of functions $r_{\max}$, $l_{\min}$, $N_{1}$ and $N_{2}$ are given in (91), (92), (93a), and (93b) in Appendix C.5, respectively. The proof of Theorem 1 is given in Appendix C.5. {Remark}[Interpretation of Theorem 1] In Theorem 1, the error tolerance $\epsilon$ and the probability $\delta$ are decomposed to provide a clearer and more structured statement. The error $\epsilon$ is broken down into four components:
  • $\epsilon_{d}$: the error caused by the noisy data;
  • $\epsilon_{l}$: the error due to the finite rollout length compared to the infinite horizon cost;
  • $\epsilon_{n}$: the error from using a finite number of rollouts in zeroth-order optimization;
  • $\epsilon_{r}$: the difference between the gradient of smoothing function $C_{r}(K)$ defined in (66) in Appendix C.2 and the original function.
The probability $\delta$ is similarly decomposed:
  • $\delta_{d}$: the concentration probabilities for gradient estimation of the smoothing function, linked to the required rollouts $N_{1}$ defined in Theorem 1;
  • $\delta_{n}$: the concentration probabilities for the noisy data, linked to the required rollouts $N_{2}$ defined in Theorem 1;
  • $\delta_{x}$: the probability that the state remains bounded.
{Theorem}
[Error bound of $\hat{\Sigma}_{K}$] Given an arbitrary tolerance $\epsilon^{\prime}>0$, which can be decomposed as $\epsilon^{\prime}=\epsilon^{\prime}_{l}+\epsilon^{\prime}_{n}+\epsilon^{\prime}_{r}$, and an arbitrary probability $\delta^{\prime}\in(0,1)$, which can be decomposed as $\delta^{\prime}=(1-\delta^{\prime}_{n})\delta^{\prime}_{x}$, for a given $K\in\mathcal{S}$, the estimated average covariance $\hat{\Sigma}_{K}$ from Algorithm 1 enjoys the following bound:
(32) $\mathbb{P}\left\{\lVert\hat{\Sigma}_{K}-\Sigma_{K}\rVert\leq\epsilon^{\prime}\right\}\geq\delta^{\prime},$
@@
(33) $r\leq r^{\prime}_{\max}(C(K),\epsilon^{\prime}_{r}),l\geq l^{\prime}_{\min}(C(K),\epsilon^{\prime}_{l}),n\geq n^{\prime}_{\min}(C(K),\epsilon^{\prime}_{r},\epsilon^{\prime}_{l},\epsilon^{\prime}_{n},\delta^{\prime}_{x},\delta^{\prime}_{n}),$
where the detailed expressions of functions $r^{\prime}_{\max}$, $l^{\prime}_{\min}$ and $n^{\prime}_{\min}$ are given in (97) (98) and (99) in C.6, respectively. The proof of Theorem 31 is given in Appendix C.6. {Remark}[Interpretation of Theorem 31] In Theorem 31, similar to Theorem 1, the error $\epsilon^{\prime}$ is decomposed into three components:
  • $\epsilon^{\prime}_{l}$: the error due to the finite rollout length compared to the infinite horizon average covariance;
  • $\epsilon^{\prime}_{n}$: the error from noise in the data;
  • $\epsilon^{\prime}_{r}$: the difference between average covariance associated with $K$ and $\hat{K}_{k}$.
The probability $\delta$ is similarly decomposed:
  • $\delta^{\prime}_{n}$: the concentration probabilities for the noisy data;
  • $\delta^{\prime}_{x}$: the probability that the state remains bounded.

### v2

@@ 3 Convergence of Model-based Policy Gradient
  • policy gradient descent: $K_{i+1}=K_{i}-\eta\nabla C(K_{i})$;
@@
where $i$ is the iteration index and $i\in\mathbb{Z}_{+}$. The Gauss-Newton method, when implemented with a step size $\eta=\frac{1}{2}$, becomes equivalent to the policy iteration algorithm, which is well-known for its quadratic convergence rate [44]. To implement the Gauss-Newton method, three terms, $(R+B^{\top}P_{K_{i}}B),\nabla C(K_{i}),\Sigma_{K_{i}}$ are necessary. In contrast, policy gradient descent only requires the gradient $\nabla C(K_{i})$, but exhibits only a linear convergence rate. The natural policy gradient method strikes a balance between the two, requiring less information than the Gauss-Newton method while providing faster convergence properties than the policy gradient descent method.
3.1 Policy gradient descent
In this subsection, we present the theorem guaranteeing the convergence of the model-based policy gradient descent method, based on the properties introduced in Sections 2.1 and 2.2.
Theorem 1 (Policy gradient descent with Adaptive Step Size)
Suppose the initial $K_{0}\in\mathcal{S}$, and consider the policy gradient descent iteration
$K_{i+1}=K_{i}-\eta_{i}\nabla C(K_{i}),\quad\forall i\in\mathbb{Z}_{+},$ (14)
where the step size satisfies $\eta_{i}\leq h_{\mathrm{PGD}}(C(K_{i}))$, and $h_{\mathrm{PGD}}(C(K_{i}))$ is defined in (62) in Appendix B.1. Then, the following relationship holds:
$C(K_{i+1})-C(K^{*})\leq\left(1-\frac{2\eta_{i}\lambda_{1}{(R)}\lambda_{1}^{2}(\Sigma_{w})}{\lVert\Sigma_{K^{*}}\rVert}\right)(C(K_{i})-C(K^{*})).$ (15)
For $\eta_{i}=h_{\mathrm{PGD}}(C(K_{i})),i\in\mathbb{Z}_{+}$ and given any accuracy gap $\epsilon>0$, if the number of iterations $N$ satisfies:
@@ 3.1 Policy gradient descent
The proof of Theorem 1 can be found in Appendix B.1. From Theorem 1, we see that the policy gradient descent method can achieve any desired accuracy $\epsilon$ within a finite number of iterations $N$ and converges linearly. According to Theorem 1, the step size is adaptive according to the cost $C(K_{i})$ at the current iterate $K_{i}$.
Remark 2 (Advantage of Adaptive Step Size over Fixed Step Size)
We can further characterize the advantage of using adaptive step sizes as proposed in Theorem 1. In [9], a fixed step size $\eta=h_{\mathrm{PGD}}(C(K_{0}))$ was proposed. While for both choices of step sizes, we can provide that $C(K_{i+1})\leq C(K_{i})$ from Theorem 1. We can also show that, given the same initial $K_{0}$ and together with the expression of $h_{\mathrm{PGD}}$, $\eta=\eta_{0}\leq\eta_{i},\forall i\in\mathbb{Z}_{+}$. This leads to the following inequality:
@@
This allows the improvement achieved with an adaptive step size to be quantified.
@@ 3.2 Natural Policy Gradient
$\theta\leftarrow\theta-\eta G_{\theta}^{-1}\nabla C(\theta),$ (16)
@@
$\hat{G}_{\bar{K}}:=\lim_{T\rightarrow+\infty}\frac{1}{T}\sum_{t=0}^{T}~\mathbb{E}\left[\nabla\mathop{log}\pi_{\bar{K}}(u_{t}|x_{t})\nabla\mathop{log}\pi_{\bar{K}}(u_{t}|x_{t})^{\top}\right].$ (18)
@@
$K_{i+1}=K_{i}-\eta\nabla C(K_{i})\Sigma_{K_{i}}^{-1}$ (19)
The explicit derivation of (19) starting from (16) and using the particular policy parametrization (17) is given in Appendix B.2. We now present the theorem that establishes the convergence guarantee of the natural policy gradient method
Theorem 2 (Natural Policy Gradient with Adaptive Step Size)
Suppose the initial $K_{0}\in\mathcal{S}$, and consider the natural policy gradient iteration
$K_{i+1}=K_{i}-\eta_{i}\nabla C(K_{i})\Sigma_{K_{i}}^{-1},\quad\forall i\in\mathbb{Z}_{+},$ (20)
@@
$\eta_{i}\leq\frac{1}{2\lVert R\rVert+\frac{2\lVert B\rVert^{2}C(K_{i})}{\lambda_{1}(\Sigma_{w})}}.$ (21)
@@
$C(K_{i+1})-C(K^{*})\leq\left(1-\frac{2\eta_{i}\lambda_{1}{(R)}\lambda_{1}(\Sigma_{w})}{\lVert\Sigma_{K^{*}}\rVert}\right)(C(K_{i})-C(K^{*})).$ (22)
@@
$\eta_{i}=\frac{1}{2\lVert R\rVert+\frac{2\lVert B\rVert^{2}C(K_{i})}{\lambda_{1}(\Sigma_{w})}},\quad\forall i\in\mathbb{Z}_{+},$ (23)
@@
The proof of Theorem 2 relies on Lemma 2 and follows the procedure outlined in [9, Lemma 15]. Theorem 2 establishes the global convergence properties of the natural policy gradient method, when an appropriate step size $\eta_{i}$ is selected. Similar to the observation in Remark 2, the introduction of adaptive step sizes improves the convergence rate, as the step sizes increase adaptively with the decreasing of the cost.
@@ 3.3 Gauss-Newton Method
We now present the theorem that provides the convergence guarantee for the Gauss-Newton method.
Theorem 3 (Gauss-Newton Method)
Suppose the initial $K_{0}\in\mathcal{S}$, and consider the Gauss-Newton iteration
$K_{i+1}=K_{i}-\eta(R+B^{\top}P_{K_{i}}B)^{-1}\nabla C(K_{i})\Sigma_{K_{i}}^{-1},\quad\forall i\in\mathbb{Z}_{+},$ (24)
@@
$C(K_{i+1})-C(K^{*})\leq\left(1-\frac{2\eta\lambda_{1}(\Sigma_{w})}{\lVert\Sigma_{K}^{*}\rVert}\right)(C(K_{i})-C(K^{*})).$ (25)
@@
The proof of Theorem 3 builds on the property shown in Lemma 2 and follows similar steps to the steps in [9, Lemma 14]. It is important to note that the choice of step size here is independent of the specific system parameters. To achieve the fastest convergence rate, one can always select the step size $\eta=\frac{1}{2}$, which is equivalent to the policy iteration algorithm.
@@ 3.4 Lack of robustness of model-based PG methods
$x_{t+1}=\underbrace{\left[\begin{array}[]{ccc}1.01&0.01&0\\ 0.01&1.01&0.01\\ 0&0.01&1.01\end{array}\right]}_{A}x_{t}+\underbrace{\left[\begin{array}[]{ccc}1&0&0\\ 0&1&0\\ 0&0&1\end{array}\right]}_{B}u_{t}+w_{t}.$ (26)
@@
$Q=0.001I_{3},\quad R=I_{3}$ (27)
The initial condition ${K}_{0}$ is chosen as the optimal gain for the LQR problem with $(A,B,50Q,R)$. To model gradient estimation errors due to noisy data, we define the gradient estimate as follows:
$\hat{\nabla}C(K):=\nabla C(K)+\Delta,$ (28)
where $\Delta\in\mathbb{R}^{3\times 3}$ is a matrix with entries $\Delta_{ij}$ sampled according to the distribution $\mathcal{N}(0,\sigma^{2})$, where the values of $\sigma\geq 0$ will be discussed later. The policy gradient descent algorithm updates the gain ${K}_{i}$ using (14) with fixed step size and the gradient estimate.
@@
When the gradient is exactly known (i.e. $\sigma=0$, black dotted line), we can use theoretical bounds on the step size to guarantee convergence of the gradient descent algorithm. In this case, the converging behavior of $\eta=0.12$ to the optimal policy is shown as an example. When the gradient estimation is affected by a small amount of noise ($\sigma=0.03$, blue dashed line), the same step size yields convergence but to a suboptimal value of gain $K$. As the noise level increases ($\sigma=0.6$, green dash-dotted line), the algorithm starts showing non-converging behavior. In this case, reducing the step size addresses the issue, as shown by the red solid line where the step size is empirically decreased to $0.01$. It is worth observing that this reduction in step size results in a slower convergence rate.
These results exemplify that the convergence guarantees obtained in the model-based scenario no longer hold, when PG algorithms are run using finite noisy data trajectories to compute gradients (or equivalently covariances). Specifically, the algorithm might converge to suboptimal solutions or even diverge. It is also shown that, in the latter case, by properly modifying the algorithms’ parameters (such as the step size), converging behaviors can be achieved.
To characterize the fundamental limitations of such algorithms and improve their performance it is then paramount to: quantify the uncertainty on gradients and covariances due to the use of finite noisy data; and analyze how this uncertainty propagates in the policy optimization algorithm. To the best of the authors’ knowledge, there is no holistic analysis that captures all these aspects and provides guarantees for model-free PG algorithms in the presence of noisy trajectory data. The main technical contribution of this work is to bridge this gap by providing: probabilistic guarantees of convergence and suboptimality bounds; closed-form expressions for the algorithm’s parameters (e.g. step size, variance reduction) that guarantee converging behavior.
@@ 4 Convergence of Model-free Policy Gradient
In this section, we analyze two variants of model-free policy gradient methods: policy gradient descent and natural policy gradient. Unlike model-based approaches, the system dynamics $(A,B)$ are unknown in the model-free setting. Instead of relying on system identification, we use zeroth-order optimization, where gradients and covariance estimates are estimated directly from trajectory data. In Section 4.1, we begin by analyzing estimation algorithms for two fundamental components utilized in the proposed methods: the cost function gradient and the covariance. We derive novel finite-sample error bounds that explicitly account for the effects of noisy data. Building on these quantitative bounds, Section 4.2 introduces a variance reduction technique to provably reduce the required number of rollouts. Leveraging these results, we establish probabilistic convergence guarantees for PGD (Section 4.3) and NPG (Section 4.4), emphasizing the more general scenario with time-adaptive step sizes. In Section 4.3 and 4.4, we denote by $\hat{K}_{i}$ the gain at the $i$-th iteration, updated using gradient estimates.
@@ 4.1 Model-free Gradient and Covariance Estimation
for k=1,,nk=1,...,n do
   1. Generate a sample gain matrix $\bar{K}_{k}=K+U_{k}$, where $U_{k}$ is drawn uniformly at random over matrices of compatible dimensions with Frobenius norm $r$;
@@ 2. Generate a sample initial state $x_{0}^{(k)}$ with $\lVert x_{0}^{(k)}\rVert\leq L_{0}$;
   3. Excite the closed-loop system with $u_{t}^{(k)}=\bar{K}_{k}x_{t}^{(k)}$ for $l$-steps starting from $x_{0}^{(k)}$, yielding the state sequence $\left\{x_{t}^{(k)}\right\}_{t=0}^{l-1}$ originating from (1b);
   4. Collect the empirical cost estimate $\hat{C}_{\bar{K}_{k}}:=\frac{1}{l}\sum_{t=0}^{l-1}x_{t}^{(k)~\top}(Q+\bar{K}^{\top}_{k}R\bar{K}_{k})x_{t}^{(k)}$ and the empirical covariance matrix $\hat{\Sigma}_{k}=\frac{1}{l}\sum_{t=0}^{l-1}~x_{t}^{(k)}x_{t}^{(k)~\top}$;
@@
return  Gradient estimate $\hat{\nabla}C(K):=\frac{1}{n}\sum^{n}_{k=1}\frac{n_{x}n_{u}}{r^{2}}\hat{C}_{\bar{K}_{k}}U_{k}$ and covariance estimate $\hat{\Sigma}_{K}:=\frac{1}{n}\sum^{n}_{k=1}\hat{\Sigma}_{k}$.
The empirical cost $\hat{C}_{\bar{K}_{k}}$ can be computed either by using the weighting matrix or directly from performance measurements, without requiring knowledge of $Q$ and $R$, as stated in Remark 1. We present two theorems to analyze the estimation errors of both $\hat{\nabla}C(K)$ and $\hat{\Sigma}_{K}$, obtained from Algorithm 1. The estimation error of $\hat{\nabla}C(K)$, denoted by $\epsilon$, together with its failure probability $\delta$, arises from several sources:
  • $\epsilon_{d}$: error caused by the noisy data;
  • $\epsilon_{l}$: error due to the finite rollout length compared to the infinite horizon cost;
  • $\epsilon_{n}$: error from using a finite number of rollouts in zeroth-order optimization;
  • $\epsilon_{r}$: difference between the gradient of the original function and the smoothing function $C_{r}(K)=\mathbb{E}_{U\sim\mathbb{B}_{r}}\left[C(K+U)U\right]$.
The probability consists of the following three components:
  • $\delta_{n}$: concentration probabilities for gradient estimation of the smoothing function, linked to $\epsilon_{n}$;
  • $\delta_{d}$: concentration probabilities for the noisy data, linked to $\epsilon_{d}$;
  • $\delta_{x}$: probability that the state is not bounded.
Having identified the different sources contributing to the estimation error and its associated failure probability, we are now in a position to state the main theorem.
Theorem 4 (Error Bound of $\hat{\nabla}C(K)$)
Given an arbitrary tolerance $\epsilon>0$, which can be expressed as $\epsilon=\epsilon_{d}+\epsilon_{l}+\epsilon_{n}+\epsilon_{r}$, and an arbitrary probability $\delta\in(0,1)$, which can be expressed as $\delta=1-(1-\delta_{d})(1-\delta_{n})(1-\delta_{x})$ with $\delta_{x},\delta_{n},\delta_{d}\in(0,1)$, for a given $K\in\mathcal{S}$, the estimated gradient $\hat{\nabla}C(K)$ from Algorithm 1 enjoys the following bound:
$\mathbb{P}\left\{\lVert\hat{\nabla}C(K)-{\nabla}C(K)\rVert\leq\epsilon\right\}\geq 1-\delta,$ (29)
@@
$r\leq r_{\max}(C(K),\epsilon_{r}),$ $\quad l\geq l_{\min}(C(K),\epsilon_{l}),$ (30a)
$n\geq\max(N_{1}(C(K),\epsilon_{r},\epsilon_{l},\epsilon_{n},\delta_{n},\delta_{d}),$ $N_{2}(C(K),\epsilon_{r},\epsilon_{l},\epsilon_{n},\epsilon_{d},\delta_{x},\delta_{n},\delta_{d})),$ (30b)
where the detailed expressions of functions $r_{\max}$, $l_{\min}$, $N_{1}$ and $N_{2}$ are given in (92), (93), (94a), and (94b) in Appendix C.5, respectively.
The proof of Theorem 4 is given in Appendix C.5. Similar to the estimation error of the gradient, the estimation error of $\Sigma_{K}$,i.e., $\epsilon^{\prime}$ and its associated failure probability $\delta^{\prime}$ also originate from different sources, which we outline below: the error $\epsilon^{\prime}$ consists of three components:
  • $\epsilon^{\prime}_{l}$: error due to the finite rollout length compared to the infinite horizon average covariance;
  • $\epsilon^{\prime}_{n}$: error from noise in the data;
  • $\epsilon^{\prime}_{r}$: difference between average covariance associated with $K$ and $\bar{K}_{k}$.
The probability $\delta$ consists of two components:
  • $\delta^{\prime}_{n}$: concentration probabilities for the noisy data;
  • $\delta^{\prime}_{x}$: probability that the state remains bounded.
Theorem 5 (Error bound of $\hat{\Sigma}_{K}$)
Given an arbitrary tolerance $\epsilon^{\prime}>0$, which can be expressed as $\epsilon^{\prime}=\epsilon^{\prime}_{l}+\epsilon^{\prime}_{n}+\epsilon^{\prime}_{r}$, and an arbitrary probability $\delta^{\prime}\in(0,1)$, which can be expressed as $\delta^{\prime}=1-(1-\delta^{\prime}_{n})(1-\delta^{\prime}_{x})$ with $\delta^{\prime}_{n},~\delta^{\prime}_{x}\in(0,1)$, for a given $K\in\mathcal{S}$, the estimated average covariance $\hat{\Sigma}_{K}$ from Algorithm 1 enjoys the following bound:
$\mathbb{P}\left\{\lVert\hat{\Sigma}_{K}-\Sigma_{K}\rVert\leq\epsilon^{\prime}\right\}\geq 1-\delta^{\prime},$ (31)
@@
$r\leq r^{\prime}_{\max}(C(K),\epsilon^{\prime}_{r}),l\geq l^{\prime}_{\min}(C(K),\epsilon^{\prime}_{l}),n\geq n^{\prime}_{\min}(C(K),\epsilon^{\prime}_{r},\epsilon^{\prime}_{l},\epsilon^{\prime}_{n},\delta^{\prime}_{x},\delta^{\prime}_{n}),$ (32)
where the detailed expressions of functions $r^{\prime}_{\max}$, $l^{\prime}_{\min}$ and $n^{\prime}_{\min}$ are given in (98) (99) and (100) in the Proof C.6, respectively.
The proof of Theorem 5 is given in Appendix C.6.